OpenStudy (31356):

The cone shown has a radius of 5 feet and a slant height of 8 feet. Which choice is closest to the total surface area? Use 3.14 to approximate pi. MEDAL REWARDED!!

OpenStudy (31356):

A) 126 B) 209 C) 204 D) 628

OpenStudy (31356):

@jigglypuff314

OpenStudy (anonymous):

pi*r^2 + pi*r*s 3.14*5^2 + 3.14*5*8 You can simplify that. Remember order of operations of course.

OpenStudy (31356):

Okay

OpenStudy (31356):

C!

OpenStudy (31356):

Can you help me with this one?

OpenStudy (31356):

Thanks!

OpenStudy (anonymous):

Sure.

OpenStudy (31356):

The cone shown has a diameter of 12 centimeters and a slant height of 8 centimeters. Which choice is closest to the total surface area? Use 3.14 to approximate pi.

OpenStudy (31356):

A)151 B)301 C)264 D)754

OpenStudy (anonymous):

Formula is pi*r^2 + pi*r*s. r=12/2=6 s=8 3.14*6^2 + 3.14*6*8

OpenStudy (31356):

Thanks!

OpenStudy (31356):

C!